OpenStudy (anonymous):

Cards are drawn with replacement from a standard shuffled deck repeatedly until a black 8 appears 1) What is the probability of the game stopping on exactly the 5 th card? 2) What is the probability of the game lasting at least 17 cards draws?

OpenStudy (shamil98):

what game

OpenStudy (anonymous):

is not a game is it just a geomtric density

OpenStudy (anonymous):

a bit complec

OpenStudy (rizags):

First one is a Bernoulli trial of the form:\[\large \left(\begin{matrix}5 \\ 1\end{matrix}\right)(\frac{1}{52})^1(\frac{51}{52})^4\times\frac{1}{5}\]

OpenStudy (anonymous):

k

OpenStudy (rizags):

The second one is extremely tedious using bernoulli trials so there has to be a better way

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

a) (50/52) ^4 * 2/52

OpenStudy (anonymous):

where did the 4 come from ?

OpenStudy (anonymous):

since we want probability of P( F F F F S ) where F = non black 8 S = black 8

OpenStudy (anonymous):

F stands for failure, S for success

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4th is success and 1 is failure

OpenStudy (anonymous):

also why is 2/52?

OpenStudy (anonymous):

i just wonder

OpenStudy (anonymous):

there are two black 8 cards

OpenStudy (anonymous):

http://www.jfitz.com/cards/classic-playing-cards.png

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

those 52 cards, 2 are black 8 i see

OpenStudy (anonymous):

i guess we still need to figure out the second one

OpenStudy (anonymous):

let X = number of cards till you drawing a black 8 you want P( X >= 17) = 1 - P( X < 17)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

let X = number of cards till you draw a black 8 you want P( X >= 17) by complement rule: P( X >= 17) = 1 - P( X <17) = 1 - P( X<= 16)

OpenStudy (rizags):

IN that case, I think the second one would be:\[\large 1-\sum_{n=0}^{15}(\frac{50}{52})^n(\frac{2}{52})\]

OpenStudy (anonymous):

does that mean n is 0 or 17?

OpenStudy (anonymous):

n is like an index n=0,1,2,3...15

OpenStudy (rizags):

that fills cases from 1 to 16, inclusive. But n is just the index of summation like @jayzdd said. The Summation can be done with a graphing calculator or online.

OpenStudy (anonymous):

that going to be a huge calculation

OpenStudy (rizags):

not for a computer

OpenStudy (anonymous):

ya i know but using wolfram is going to take me while to find the command

OpenStudy (anonymous):

wolfram is user friendly, you dont need to know the command. i have used trial and error. a few times i have had to look up a command, for sophisticated problems

OpenStudy (anonymous):

alright thank you for showing me the command on wolfram

OpenStudy (anonymous):

i just google it for command lol

OpenStudy (anonymous):

no thats not the command

OpenStudy (anonymous):

we are using plain english

OpenStudy (anonymous):

btw is 16 not 15

OpenStudy (anonymous):

wolfram uses mathematica language, which has special syntax or it can use english if its precise enough to understand it

OpenStudy (anonymous):

but anywhow thank so much for clarification

OpenStudy (rizags):

its 15 because I start at 0 for a specific reason

OpenStudy (anonymous):

isn't the "lasting at least 17 cards"

OpenStudy (rizags):

at least means greater than or equal to 17, so I set up the sum to start at 0 and head up to 15, which functionally translates as 16 because of the corresponding Bernoulli Trial

OpenStudy (anonymous):

that make more sense thank you